In outer space rock 1, with mass 4 kg and velocity < 4300, -2600, 2000 > m/s, struck rock 2, which was at rest. After the colMuch appreciated!
Answer #1

Rock 1 ма Initimassam, = 4 kg a velocity w ni = (43007 -26003+2000 Å 1 m/s Final velocity a Vgs = (42001 - 2000Î +2400) m/s I= 900f - 2400-1600R bet = <800, -2400,-1600 kg.mis. । २m. kinetic energy YN (17200)?+(-104001? + 1800072 } 2x4 Kai = 5.85 x 1If, AE thermal at AE thermal, 2 = 2.68x1065 Then Kas = (kait kai) - Kef - 2.68*106 = 536x109 - 268x104 268x104g 1 K2f = 2,68x